cot2 theta + 5 = 6cot theta. Present your answer in radians (where possible in exact form, otherwise – up to three decimal places).
Answer by lwsshak3(11628)

cot2 theta + 5 = 6cot theta. Present your answer in radians (where possible in exact form, otherwise – up to three decimal places)
**
cot^2x+5=6cotx
cot^2x-6cotx+5=0
(cotx-5)(cotx-1)=0
cotx=5
Use inverse tan keys on calculator
tanx=1/5
x≈.197
or
cotx=1
tanx=1
x=π/4